OpenStudy (anonymous):

2y + 7x = 4; x= 5, 10, 15 SHOW WORK PLEASE

OpenStudy (anonymous):

x=5 2y+7(5)=4 2y+35=4 2y=-31 y=-31/2 x=10 2y+7(10)=4 2y+70=4 2y=-66 y=-33 x=15 2y+7(15)=4 2y+105=4 2y=-101 y=-101/2